OpenStudy (anonymous):

hey i need help plzzzz -7/(x^2-2x-15)

OpenStudy (anonymous):

you trying to solve? graph?

OpenStudy (anonymous):

What is the excluded value?

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

We need to find where x^2-2x-15=0 now we can use the quadratic formula to find x x=-b+/-sqrt(b^2-4ac)/2a here a=1 b=-2 and c= -15 x=2+/-sqrt(-2^2-4(1)(-15)/2(1) x=2+/-sqrt(64)/2 x=2+/-8/2 so x=2+8/2=5 or x=2-8/2=-3 the excluded values* are 5 and -3

OpenStudy (anonymous):

sorry im bad at latex

OpenStudy (anonymous):

hey ty

OpenStudy (anonymous):

that make sense right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can check by plugging them back in the denominator=0 so it undefined at those values